OpenStudy (anonymous):

Will give medal and become a fan! Simplify. square root of 3 times square root of 21 square root of 24 square root of 63 3 square root of 7 None of the above

OpenStudy (anonymous):

\[\sqrt{3} * \sqrt{21}\]

OpenStudy (anonymous):

3 square root of seven

OpenStudy (anonymous):

Square root of 21 is 4.58257695 Square root of 3 is 1.732050808.

OpenStudy (anonymous):

Multiply them and square the answer.

OpenStudy (anonymous):

So it would be 3 square root of seven ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What about this one? Simplify:square root of 5 open parentheses 8 plus 3 square root of 6 close parentheses

OpenStudy (anonymous):

\[\sqrt{3} \times \sqrt{21} = \sqrt{63} = \sqrt{3 \times 3 \times 7} = 3\sqrt{7}\] @wildchild588 : Your question is answered with the working

OpenStudy (anonymous):

\[\sqrt{5}*(8+[3*\sqrt{6}])\]

OpenStudy (cj49):

√(3) * √(21) = √(3 * 21) = √(63)

OpenStudy (anonymous):

The brackets are just another version of parenthesis, but I'm only trying to say to do 3+sqrt of 6 before anything else.

OpenStudy (anonymous):

These are the possible outcomes. 8 square root of 5 end root plus 3 square root of 30 5 square root of 10 130 None of the above

OpenStudy (anonymous):

Simplify the equation I gave you.

OpenStudy (anonymous):

So it would be 5 square root of 10